OpenStudy (anonymous):

(a-b)^2 over b-a simplify

OpenStudy (anonymous):

The trick is to factor out a -1 from the (b - a). Notice that if you do this, you get -(a - b) so we have (a-b)(a-b) / -(a-b) cancel one term leaves us with (a-b) / -1 = -(a-b) = b-a

OpenStudy (anonymous):

ty soooo much [; garettm

OpenStudy (anonymous):

np